Angad is 10 years older than Alex. Mark is 11 years younger than Alex. If the total of their ages is 68, how old is the youngest of them?
Found 2 solutions by mananth, MathTherapy:
Answer by mananth(16946)

Angad is 10 years older than Alex. Mark is 11 years younger than Alex. If the total of their ages is 68, how old is the youngest of them?
Let Alex's age be x
Angad's age will be x+10
Add their ages
Mark's age will be x-11
x+x+10+x-11 =68
3x-1=68
3x=69
x=69/3 = 23
Angad will be 23+10 =33
Mark's age will be 23-11 =12

Answer by MathTherapy(10552)


Angad is 10 years older than Alex. Mark is 11 years younger than Alex. If the total of their ages is 68, how old is the youngest of them?
After reading, it's quite obvious that Mark is the youngest 
Let Mark's age be M
Then Alex' is: M + 11, and Angad's is: M + 11 + 10 = M + 21
With all ages summing to 68, we have: M + M + 11 + M + 21 = 68
3M = 36
The youngest, who's Mark, or